Find the curve in the xy-plane that passes through the point each point is sec23x. nd whose slope at
Answer #1

Answer

Let the curve in the xy-plane be

y = f(x)

Since the slope at each point of the curve is sec2(3x),

\frac{\mathrm{d} y}{\mathrm{d} x}=\sec^2(3x)

By integration, we will get

y=\int\sec^2(3x)\, dx =\frac{\tan(3x)}{3}+C

where, C is the constant of integration.

Hence, the equation of the curve is

y =\frac{\tan(3x)}{3}+C

Since the curve passes through the point (\pi/12, 2),

2 =\frac{\tan\left ( 3\times \frac{\pi}{12} \right )}{3}+C

\Rightarrow 2 =\frac{\tan\left ( \frac{\pi}{4} \right )}{3}+C

\Rightarrow 2 =\frac{1}{3}+C

\Rightarrow C = 2-\frac{1}{3}=\frac{5}{3}

Therefore, the equation is

{\color{Blue} y =\frac{\tan(3x)}{3}+\frac{5}{3}}